OpenStudy (sarahezzmck):

Help Please! What statement correctly describes the key features of the graph of f(x) = − 4(one half)x + 1 − 3?

OpenStudy (sarahezzmck):

\[f(x)=-4\left(\begin{matrix}1 \\ 2\end{matrix}\right)^{x+1}-3\]

OpenStudy (sarahezzmck):

Y-intercept of (0, −5), starts up on the left, gets closer to y = −3 on the right Y-intercept of (0, −5), starts down on the left, gets closer to y = −3 on the right Y-intercept of (0, 5), starts up on the left, gets closer to y = −3 on the right Y-intercept of (0, 5), starts down on the left, gets closer to y = −3 on the right

Parth (parthkohli):

Let's start with the y-intercept. What is \(f(x)\) when \(x = 0\)?

OpenStudy (sarahezzmck):

it is f(0)=0 ?

Parth (parthkohli):

Nope. Plug \(x = 0\) in that equation you are given.

OpenStudy (sarahezzmck):

im not sure how to do that i forgot how :( can u show me how to do that?

Parth (parthkohli):

\[f(x) = -4 \left(\dfrac{1}{2}\right)^{x + 1} - 3\]\[f(0) = -4\left(\dfrac{1}{2}\right)^{0+1} - 3\]

OpenStudy (sarahezzmck):

oh, ok, so when you do that it just changes the x value in the x+1 so it would be o and then the f of x is zero as well?

Parth (parthkohli):

Why would \(x+1\) be zero?

OpenStudy (anonymous):

IDK

OpenStudy (sarahezzmck):

because x=0 so the x's in the equation change to zero?

OpenStudy (sarahezzmck):

shhhh Alex and turn ur caps off jeez lol

OpenStudy (anonymous):

ok

Parth (parthkohli):

The only \(x\) there is \(x + 1\). When you plug in zero, you get \(1\).

OpenStudy (sarahezzmck):

oh ok, because 0+1 =1

Parth (parthkohli):

Hey, Rolltideroll!

Parth (parthkohli):

Right. So what is\[-4\left(\dfrac{1}{2}\right)^1 -3\]

OpenStudy (anonymous):

Hey whats up

Parth (parthkohli):

Not much! How're you doing? @SarahEZZMcK Could you find it?

OpenStudy (anonymous):

doing good have we meet before

OpenStudy (sarahezzmck):

i have no clue im really sorry! im terrible at math

OpenStudy (anonymous):

I dont understand this at all

OpenStudy (sarahezzmck):

@razor99 @Warriorslayer @adilalvi @Awesome781 @bibby @Darry&Gator @FlvsGirl help please!!!!

OpenStudy (anonymous):

what about the thing you used the other night

OpenStudy (sarahezzmck):

dont have it now

OpenStudy (anonymous):

YYYYYY

OpenStudy (sarahezzmck):

bc im a loser why else?

OpenStudy (sarahezzmck):

@ztp5031

OpenStudy (sarahezzmck):

???

OpenStudy (anonymous):

no you are not a loser

OpenStudy (anonymous):

any one that is my friend is not a loser

OpenStudy (sarahezzmck):

@DK94 help please

OpenStudy (anonymous):

@chucky00 please help Sarah

OpenStudy (sarahezzmck):

pleaaaaasseeeeeeeeeee

OpenStudy (anonymous):

try this link http://openstudy.com/study#/updates/529970bae4b0939d132bab75

OpenStudy (adilalvi):

sorry

OpenStudy (sarahezzmck):

its ok, and i tried that answer Alex, it was wrong :(

OpenStudy (anonymous):

oh ok sorry i coulnt help:((((

OpenStudy (sarahezzmck):

@BeginnerCountrySinger Elijah help! @djcool31

OpenStudy (sarahezzmck):

its ok Alex :)

OpenStudy (beginnercountrysinger):

HOW IN THE.... I CANT DO THIS :3

OpenStudy (sarahezzmck):

ugh ok well igtg

OpenStudy (beginnercountrysinger):

byyyyyyeeeeee

OpenStudy (sarahezzmck):

Darry! Can u help????

OpenStudy (anonymous):

I suck at math. I'll see what i can do

OpenStudy (sarahezzmck):

ok thanks! :)

OpenStudy (anonymous):

Yeah i don't know. I am so confused

OpenStudy (sarahezzmck):

@***[ISURU]*** can u help?? And yeah me too! I loooked it up, and it said the answer was C, but it was the wrong answer

OpenStudy (anonymous):

@tHe_FiZiCx99 @agent0smith @mathmale @tkhunny @ranga @CGGURUMANJUNATH

OpenStudy (sarahezzmck):

@***[ISURU]*** can u help?? And yeah me too! I loooked it up, and it said the answer was C, but it was the wrong answer

OpenStudy (anonymous):

ok i guessed this and i got it right on flvs this is the correct answer, lol dont ask me how but i got it right Y-intercept of (0, −5), starts up on the left, gets closer to y = −3 on the right so b?